A manufacturing company leases a building for $90000 per year for its manufacturing facilities. In addition, the machinery in this building is being paid for in installments of $22000 per year. Each unit of the product produced costs $14 in labor and $9 in materials. The product can be sold for $45

How many units per year must be sold for the company to breakeven? Choose the closest ans.

A. 8146

B. 6109

C. 7127

D. 3055

E. 5091
Answer #1

Let the output be Q

Fixed cost = lease + instalment on machines = 90000 + 22000 = 112000

Selling price per unit = 45

Cost price per unit = material cost + labor cost=14+9 = 23

Now Total cost = 23*Q + 112000

Total revenue = 45*Q

At breakeven

Total revenue = total cost

45Q=23Q+112000

22Q =112000

Q=5090.90 = 5091

Answer is option E